A large multinational shipping company, MGC, Inc. has just decided to spend €10 million on new storage space in Munich, €45 million on new aircraft, and €5 million on additional acquisition of kerosene. In addition to these expenses, the company is producing 5 million parcels at a price of €5 per parcel. Now suppose that MGC plans to have a tenth of that production in inventory. Over time, the company’s parcels have met with increasing demand, but the inventory has only increased by €1 million.

a. What is this firm’s total planned investment?

b. How much did the firm actually invest?

c. What is the difference between actual and planned investment? Should MGC produce more or fewer parcels? Why?



1
Expert's answer
2022-04-06T06:57:15-0400

A. Investment is an asset bought not for current consumption, so we don't include kerosene and nine tenth of parcels. So, planned total consumption is 10+45+0.1×25=57.5 million

B. If the inventory increased only by 1 mln, thus the firm invested 10+45+1=56 mln

C. The difference is 57.5-56=1.5 mln, the firm should produce more parcels to cover the shortage